Static Friction

A cyclist competes in a one-lap race around a flat, circular course of radius 140m . Starting from rest and speeding up at a constant rate throughout the race, thecyclist covers the entire course in 60s. The mass of the bicycle (including the rider) is 76kg. The net force acting on the bicyclist at the finish line is 470 N. Whatis the minimum coefficient of static friction between the race track and the bicycle?
Answer #1
F=uN=470
u=470/N=470/(76*9.81)=.630